The mass of a hot-air balloon and its cargo (not including the air inside) is 200 kg. The air outside is at 10.0°C and 101 kPa. The volume of the balloon is475 m3. To what temperature must the air in the balloon be heated before the balloon will liftoff? (Air density at 10.0°C is 1.25 kg/m3.)
1 K
Answer #1
Total weight of the balloon + cargo system + filled airis
V*D*g + 200*9.8 = 475*D*9.8 + 1960 N, where D is thedensity of the air filled nside and V is the volume of theballoon.
Buoyant force acting on the balloon is d*V*g,
where d is the density of the air outside, V is volumeof the balloon,
So, for the balloon to lift off,
Buoyant force must be greater than the totalforce.
So, for d = 1.25 kg/m3,
V = 475 m3
1960 + 475*9.8*D = 1.25*475*9.8
So, D = 0.82895 kg/m3
So, using PV/nT = constant,
n/V α P/T
(mn/V) α P/T , where m is the average molarmass.
So, D α P/T
Thus, D/d = t/T,
t = 10 + 273 K = 283 K,
D = 0.82895 kg/m3, d = 1.25 kg/m3
T = 283*(1.25/0.82895) K = 426.746 K